TOO SHORT CN BOND LENGTHS EXPERIMENTALLY FOUND IN COBALT CYANIDE: AN AB INITIO MOLECULAR ORBITAL STUDY

\underline{REI OKUDA}, TSUNEO HIRANO, and UMPEI NAGASHIMA, Grid Technology Research Center, Institute of Advanced Industrial Science and Technology, 6-9-3 Ueno, Taito-ku, Tokyo 110-0015, Japan.

In the previous Ohio meeting, we pointed out that the CN bond lengths experimentally found for FeNC, CoCN, and NiCN are too-short.^{a} The CN bond lengths in these radicals found by spectroscopy are shorter in this order than those predicted by high-level ab initio molecular orbital calculations. The tendency is in parallel with the expected ionicity for the metal-N or metal-C bond, and hence is in parallel with the floppiness in bending motion. Recently submillimeter spectra of X̃^{3}Φ_{i} has been published by Sheridan, Flory, and Ziurys, and the CN bond length r_{0} derived for the X̃^{3}Φ_{Ω=4} is reported to be 1.1313(10) Å.^{b} Our r_{e} value for the CN bond predicted at the level of the MR-SDCI+Q+Relativistic-correction/Roos ANO(Co, C, N) is 1.171 Å, which is in the normal range observed and predicted for many CN-containing molecules. The difference in r_{0} and r_{e} shows how floppy CoCN is for the bending vibration mode.

aT. Hirano, R. Fukui, and U. Nagashima, 59th Ohio State Univ. Internat. Sympo. Mol. Spectrosc., RF05 (2004).

bP. M. Sheridan, M. A. Flory, and L. M. Ziurys, J. Chem. Phys., 121, 8360-8368 (2004).